# Robust Measurement of Stellar Streams Around the Milky Way: Correcting Spatially Variable Observational Selection Effects in Optical Imaging Surveys

### v2

@@ 4.3 DES Corrections
Figure 5: The effective weight map (corrected / uncorrected), smoothed with a $0.15^{\circ}$ kernel, for stellar objects in our faintest magnitude bin. This map is used for the validation tests in Section 5.
@@
Before applying these corrections to the Phoenix stellar stream specifically, we validate our methodology (Section 5). For our tests, we construct an effective weight map for stellar objects by taking corrected divided by original counts in Figure 3. The smoothed effective weight map is shown in Figure 5 with the boundary removed as it suffers from additional systematics. Bins of this map are used to sample a large range of impacts on stars from survey properties.
5 Validation of Methodology
This section tests the accuracy of the probabilities calculated in Section 3. We test the total corrective power of our relative detection rates (Section 5.1), how this changes with the number of objects used to train (Section 5.2), and how consistent our final corrections are as a function of the number of objects used for training (Section 5.3).
5.1 Overall Corrective Power
Figure 6: Detection rates relative to the average for a 20% testing subset of the Balrog objects before and after corrections. Before corrections is shown in blue and after corrections in orange. The plot titles describe which relative detection rates are being shown. A drop in variance like observed shows that selection effects are being mitigated.
@@ 5.1 Overall Corrective Power
Subsets of the data are created by binning the healpixels based on their effective weight from Figure 5, to obtain 10 bins of equal sized samples. From each area bin, we select 80% of the Balrog objects to be in our training sample which is used to calculate corrections applied to the remaining test set. Figure 6 shows relative detection rates on the test set across the effective weight bins. Comparing rates before (blue) and after (orange) corrections shows that variations in relative detection rates for both stars and galaxies drop by a factor of $\sim 5$ .
5.2 Convergence from Object Counts
@@
Characterizing how corrective power scales with more SSI objects will help inform SSI strategies for future analyses. To investigate this, we perform the test from Section 5.1 while varying the size of the training sample. Due to noise concerns, we reduce the number of effective weight bins to 3. Our results are shown in Figure 7 where we plot the detection rates for correctly classified stars (left) and incorrectly classified galaxies (right) as a function of the size of the training dataset. Blue (orange, green) points represent an effective weight bin that started with above (approximately, below) average relative detection rates. To illustrate the general behavior we include an exponential decay envelope in red. This line shows convergence is reached with $\sim$40% of the training set in both cases. Beyond this point we are limited by the Poisson noise of the test dataset and cannot probe further convergence.
5.3 Repeatability of Results
@@
With Balrog Y6 covering the entire DES footprint instead of just 20% (Anbajagane et al., 2025a), we can use our fit to predict the consistency of results when $\sim 5$ times more objects are injected. In this case, we expect two full runs of Balrog Y6 to have stellar effective weight map residuals with a standard deviation of $\sim 0.03$, compared to $\sim 0.06$ for Balrog Y3. Other differences in Balrog Y6 mentioned in Everett et al. (2022) and Tabbutt (2023) such as no longer using delta stars will impact the accuracy of this prediction.
Residuals can also be used to validate our cycle limit for convergence as mentioned in Section 3. We compare our calculated effective weight map to an effective weight map where we put no limit on the number of cycles in the iterative training. The standard deviation for the residuals of these maps was 0.003, showing that the cycle limit we have chosen has allowed results to converge.
6 Application to Stellar Streams
This section applies the corrections from Section 4 to observations of stellar streams. First we look at the Phoenix stellar stream and show that our corrections cause statistically significant linear density shifts (6.1). We then look at synthetic streams and show that our correction algorithm properly suppresses artificial power at length scales larger than the DECam field of view (FOV) (6.2) of 3 sq. deg. This suppression of power is detectable on individual stream realizations at length scales relevant for subhalo interactions (Banik et al., 2021).
6.1 Phoenix Stellar Stream
Figure 9: Phoenix stellar stream before and after corrections are applied. A gaussian smoothing with a kernel of $0.15^{\circ}$ is used to smooth these maps. The (RA, Dec.) endpoints in degrees of this stream are (20.1, -55.3) and (27.9, -42.7) (Shipp et al., 2018).
@@ 6.1 Phoenix Stellar Stream
Taking the matched-filter map from Section 4.1 we show the original (top) and corrected (bottom) counts for the on-sky region around the Phoenix stream in Figure 9. It is immediately obvious that the depth feature along RA $\sim 27^{\circ}$ is removed in the corrected map, enhancing the signal of the stellar stream. For a quantitative measurement of this effect, we compute the number of excess stars along the stream relative to the background.
@@
6.2 Synthetic Stellar Streams
@@
To probe the effect of our corrections on the 1D density auto-correlation function, we compute this power spectra for 5000 simulated streams before and after corrections. These simulated streams are injected into an arbitrary region of the sky at high Galactic latitude with large variations in effective weight (the area is fixed between realizations). For each realization, we assume a uniform distribution for the true number of stars and galaxies plus a stream population of stars. Since we don’t know the absolute detection rate, we fix the number of detected objects for each realization. We then use relative detection rates to create realistic spatial distributions for all of our mocks. Our stream is defined to, on average, have a 25% excess of stars compared to the background stellar population. Corrections are then performed on the data based on the relative detection rates and correct classification probabilities as prescribed in Section 3. For more details, refer to Appendix D.
To compare these uncorrected and corrected results to a baseline, we independently do a run with perfect classification and no variations in detection rates. Two runs were performed, one with a uniform density stream, and one with a sinusoidal density stream with period $4.4^{\circ}$, two times the DECam FOV. In Figure 11 we show the results of this exercise. The lines are the median power spectra for baseline (red), corrected (orange), and uncorrected (blue) results. Shaded regions represent the $16$th and $84$th percentiles among the 5000 simulations for each point along the power spectra. For the uniform density stream (left), the corrected stream obtains a near constant power spectra while the uncorrected stream obtains extra power at scales larger than the DECam FOV. For the sinusoidal density stream (right), the corrected stream does gain signal power at the expected scales, although its amplitude is lower due to the contamination of galaxies. The differences from the uncorrected stream can be seen even on individual runs, and as is shown in Banik et al. (2021), these are the same angular scales expected to be sensitive to dark matter subhalo interactions. Dashed lines in Figure 11 represent results with a more conservative magnitude cut of $r<23.9$. Noise levels increase as expected, and in the sinusoidal case the signal to noise ratio of the peak is reduced by a factor of $10\%$ compared to the full $r<24.5$ data, showing the advantage of using fainter objects.
@@ 7 Conclusion
Section 4 shows our corrections applied to DES DR1 data. After these corrections were applied, galaxy counts became more uniform on large scales. Features correlated with survey properties such as the depth feature at RA $\sim 30^{\circ}$ and edge artifacts are also visually diminished after corrections.
Validation of our algorithm is performed in Section 5. For overall corrective power, we find that key relative detection rates in our faintest magnitude bin had their standard deviations drop by a factor of at least five after corrections. Running this same test with a variable training set size showed that relative detection rates improved in uniformity when more training objects were used, but only up to a limit. Finally, we find that the repeatability of our results improved as more training objects were used. These improvements persisted over the entire range of training set sizes that we were able to test.
We apply our corrections to stellar streams in Section 6. Here we find that the changes in the linear density of Phoenix due to our corrections are statistically significant, which calls into question some of the apparent density variations reported in the literature. For uniform density simulated streams, our corrections mitigated all signals from survey property variations. For sinusoidal density simulated streams, our corrections saw a signal at the proper frequency, although the amplitude of the corrected signal was reduced.
One clear next step is to derive corrections for the DES Y6 run of Balrog (Anbajagane et al., 2025a). This SSI run has injections across the entire DES footprint compared to the 20% coverage for Balrog Y3. One large advantage of this will come in the form of testing set size for the tests performed in Section 5. The number of sources in a 20% testing set of the Balrog Y6 data would be on the order of the entire Balrog Y3 catalog. This will make Balrog Y6 less susceptible to potential bias in estimates of relative detection rates that are more likely to arise with smaller sample sizes. This could lead to the relative detection rates in Figure 7 moving closer to one in the future if the current limits on accuracy were due to biased testing sets. Additionally, this framework could easily be applied to data from the DECam Local Volume Exploration survey (DELVE; Drlica-Wagner et al. 2021) processed through the DECam All Data Everywhere (DECADE; Anbajagane et al. 2025b) campaign. This data set uses the same instrument with the same data reduction pipeline, but contains much more heterogeneous survey properties.
@@ A.2 Relative Detection Rates
The calculation of these rates is performed in a similar manner as in Section 5.3.1 of Rodríguez-Monroy et al. (2022). Within a bin from some survey property, we look at the average number of detections of interest (Balrog delta stars classified as stars within our magnitude bin) per healpixel in that bin to find a characteristic detection density. This is then compared to the average number of detections of interest per healpixel across all of the valid healpixels. If the distribution of objects that could cause a detection of interest is uniform on large scales, the ratio of these two densities will then represent the rate at which detections of interest occur in the survey property bin relative to the rate at which detections of interest occur overall. Since Balrog objects are injected on a uniformly spaced hex grid and the distribution of objects on this grid is random, this approach is valid for our calculations.
@@ Appendix C Relative Detection Rate Correlations
Overall, the scatter in these relations eliminates the possibility of certain simplifications that could have potentially been made to our correction algorithm. Referring back to Section 3, our current correction algorithm, originally given in Eq. (5), is given by:
@@ Appendix D Algorithm Design Tests
In Section 3, we presented two different correction algorithms: one to estimate what the number of objects classified as a stars would have been with uniform survey properties (we will refer to it as the recovered classified algorithm), and one to estimate what the true number of observed stars would have been with uniform survey properties (recovered observed). The recovered classified algorithm was chosen for use throughout this work. This section will discuss tests performed on the two algorithms that were used in the making of that decision.

### v3

@@ IV.3 DES Corrections
Figure 5: The effective weight map (corrected / uncorrected), smoothed with a $0.15^{\circ}$ kernel, for stellar objects in our faintest magnitude bin. This map is used for the validation tests in Section V.
@@
Before applying these corrections to the Phoenix stellar stream specifically, we validate our methodology (Section V). For our tests, we construct an effective weight map for stellar objects by taking corrected divided by original counts in Figure 3. The smoothed effective weight map is shown in Figure 5 with the boundary removed as it suffers from additional systematics. Bins of this map are used to sample a large range of impacts on stars from survey properties.
V Validation of Methodology
This section tests the accuracy of the probabilities calculated in Section III. We test the total corrective power of our relative detection rates (Section V.1), how this changes with the number of objects used to train (Section V.2), and how consistent our final corrections are as a function of the number of objects used for training (Section V.3).
V.1 Overall Corrective Power
Figure 6: Detection rates relative to the average for a 20% testing subset of the Balrog objects before and after corrections by effective weight bin number. Before corrections is shown in blue and after corrections in orange. The plot titles describe which relative detection rates are being shown. A drop in variance like observed shows that selection effects are being mitigated.
@@ V.1 Overall Corrective Power
Subsets of the data are created by binning the healpixels based on their effective weight from Figure 5, to obtain $10$ bins of equal sized samples. From each area bin, we select $80\%$ of the Balrog objects to be in our training sample which is used to calculate corrections applied to the remaining test set. Figure 6 shows relative detection rates on the test set across the $10$ effective weight bins, which are referred to in the figure as the DES footprint area bins. Comparing rates before (blue) and after (orange) corrections shows that variations in relative detection rates for both stars and galaxies drop by a factor of $\sim 5$ .
V.2 Convergence from Object Counts
@@
Characterizing how corrective power scales with more SSI objects will help inform SSI strategies for future analyses. To investigate this, we perform the test from Section V.1 while varying the size of the training sample. Due to noise concerns, we reduce the number of effective weight bins to 3. Our results are shown in Figure 7 where we plot the detection rates for correctly classified stars (left) and incorrectly classified galaxies (right) as a function of the size of the training dataset. Blue (orange, green) points represent an effective weight bin that started with above (approximately, below) average relative detection rates. To illustrate the general behavior we include an exponential decay envelope in red. This line shows convergence is reached with $\sim$40% of the training set in both cases. Beyond this point we are limited by the Poisson noise of the test dataset and cannot probe further convergence.
V.3 Repeatability of Results
@@
With Balrog Y6 covering the entire DES footprint instead of just 20% (Anbajagane et al., 2025b), we can use our fit to predict the consistency of results when $\sim 5$ times more objects are injected. In this case, we expect two full runs of Balrog Y6 to have stellar effective weight map residuals with a standard deviation of $\sim 0.03$, compared to $\sim 0.06$ for Balrog Y3. Other differences in Balrog Y6 mentioned in Everett et al. (2022) and Tabbutt (2023) such as no longer using delta stars will impact the accuracy of this prediction.
Residuals can also be used to validate our cycle limit for convergence as mentioned in Section III. We compare our calculated effective weight map to an effective weight map where we put no limit on the number of cycles in the iterative training. The standard deviation for the residuals of these maps was 0.003, showing that the cycle limit we have chosen has allowed results to converge.
VI Application to Stellar Streams
This section applies the corrections from Section IV to observations of stellar streams. First we look at the Phoenix stellar stream and show that our corrections cause statistically significant linear density shifts (VI.1). We then look at synthetic streams and show that our correction algorithm properly suppresses artificial power at length scales larger than the DECam field of view (FOV) (VI.2) of 3 sq. deg. This suppression of power is detectable on individual stream realizations at length scales relevant for subhalo interactions (Banik et al., 2021a).
VI.1 Phoenix Stellar Stream
Figure 9: Phoenix stellar stream before and after corrections are applied. A gaussian smoothing with a kernel of $0.15^{\circ}$ is used to smooth these maps. The (RA, Dec.) endpoints in degrees of this stream are (20.1, -55.3) and (27.9, -42.7) (Shipp et al., 2018b).
@@ VI.1 Phoenix Stellar Stream
Using the color and magnitude cuts from Section IV.1, we show the original (top) and corrected (bottom) counts for the on-sky region around the Phoenix stream in Figure 9. It is immediately obvious that the depth feature along RA $\sim 27^{\circ}$ is removed in the corrected map, enhancing the signal of the stellar stream. For a quantitative measurement of this effect, we compute the number of excess stars along the stream relative to the background.
@@
VI.2 Synthetic Stellar Streams
@@
To probe the effect of our corrections on the 1D density auto-correlation function, we compute this power spectra for 5000 simulated streams before and after corrections. These simulated streams are injected into an arbitrary region of the sky at high Galactic latitude with large variations in effective weight (the area is fixed between realizations). For each realization, we assume a uniform distribution for the true number of stars and galaxies plus a stream population of stars. Since we don’t know the absolute detection rate, we fix the number of detected objects for each realization. We then use relative detection rates to create realistic spatial distributions for all of our mocks. Our stream is defined to, on average, have a 25% excess of stars compared to the background stellar population. Corrections are then performed on the data based on the relative detection rates and correct classification probabilities as prescribed in Section III. For more details, refer to Appendix D.
To compare these uncorrected and corrected results to a baseline, we independently do a run with perfect classification and no variations in detection rates. Two runs were performed, one with a uniform density stream, and one with a sinusoidal density stream with period $4.4^{\circ}$, two times the DECam FOV. In Figure 11 we show the results of this exercise. The lines are the median power spectra for baseline (red), corrected (orange), and uncorrected (blue) results. Shaded regions represent the $16$th and $84$th percentiles among the 5000 simulations for each point along the power spectra. For the uniform density stream (left), the corrected stream obtains a near constant power spectra while the uncorrected stream obtains extra power at scales larger than the DECam FOV. For the sinusoidal density stream (right), the corrected stream does gain signal power at the expected scales, although its amplitude is lower due to the contamination of galaxies. The differences from the uncorrected stream can be seen even on individual runs, and as is shown in Banik et al. (2021a), these are the same angular scales expected to be sensitive to dark matter subhalo interactions. Dashed lines in Figure 11 represent results with a more conservative magnitude cut of $r<23.9$. Noise levels increase as expected, and in the sinusoidal case the signal to noise ratio of the peak is reduced by a factor of $10\%$ compared to the full $r<24.5$ data, showing the advantage of using fainter objects.
@@ VII Conclusion
Section IV shows our corrections applied to DES DR1 data. After these corrections were applied, galaxy counts became more uniform on large scales. Features correlated with survey properties such as the depth feature at RA $\sim 30^{\circ}$ and edge artifacts are also visually diminished after corrections.
Validation of our algorithm is performed in Section V. For overall corrective power, we find that key relative detection rates in our faintest magnitude bin had their standard deviations drop by a factor of at least five after corrections. Running this same test with a variable training set size showed that relative detection rates improved in uniformity when more training objects were used, but only up to a limit. Finally, we find that the repeatability of our results improved as more training objects were used. These improvements persisted over the entire range of training set sizes that we were able to test.
We apply our corrections to stellar streams in Section VI. Here we find that the changes in the linear density of Phoenix due to our corrections are statistically significant, which calls into question some of the apparent density variations reported in the literature. For uniform density simulated streams, our corrections mitigated all signals from survey property variations. For sinusoidal density simulated streams, our corrections saw a signal at the proper frequency, although the amplitude of the corrected signal was reduced.
One clear next step is to derive corrections for the DES Y6 run of Balrog (Anbajagane et al., 2025b). This SSI run has injections across the entire DES footprint compared to the 20% coverage for Balrog Y3. One large advantage of this will come in the form of testing set size for the tests performed in Section V. The number of sources in a 20% testing set of the Balrog Y6 data would be on the order of the entire Balrog Y3 catalog. This will make Balrog Y6 less susceptible to potential bias in estimates of relative detection rates that are more likely to arise with smaller sample sizes. This could lead to the relative detection rates in Figure 7 moving closer to one in the future if the current limits on accuracy were due to biased testing sets. Additionally, this framework could easily be applied to data from the DECam Local Volume Exploration survey (DELVE; Drlica-Wagner et al. 2021) processed through the DECam All Data Everywhere (DECADE; Anbajagane et al. 2025a) campaign. This data set uses the same instrument with the same data reduction pipeline, but contains much more heterogeneous survey properties.
@@ A.2 Relative Detection Rates
The calculation of these rates is performed in a similar manner as in Section 5.3.1 of Rodríguez-Monroy et al. (2022a). Within a bin from some survey property, we look at the average number of detections of interest (Balrog delta stars classified as stars within our magnitude bin) per healpixel in that bin to find a characteristic detection density. This is then compared to the average number of detections of interest per healpixel across all of the valid healpixels. If the distribution of objects that could cause a detection of interest is uniform on large scales, the ratio of these two densities will then represent the rate at which detections of interest occur in the survey property bin relative to the rate at which detections of interest occur overall. Since Balrog objects are injected on a uniformly spaced hex grid and the distribution of objects on this grid is random, this approach is valid for our calculations.
@@ Appendix C Relative Detection Rate Correlations
Overall, the scatter in these relations eliminates the possibility of certain simplifications that could have potentially been made to our correction algorithm. Referring back to Section III, our current correction algorithm, originally given in Eq. (6), is given by:
@@ Appendix D Algorithm Design Tests
In Section III, we presented two different correction algorithms: one to estimate what the number of objects classified as a stars would have been with uniform survey properties (we will refer to it as the recovered classified algorithm), and one to estimate what the true number of observed stars would have been with uniform survey properties (recovered observed). The recovered classified algorithm was chosen for use throughout this work. This section will discuss tests performed on the two algorithms that were used in the making of that decision.